OpenStudy (anonymous):

Find the slope. x: -2, -1,0,1 y: 2, 1, 0, 1 Select one: a. 1 b. -1 c. 1/2 d. -1/2

OpenStudy (anonymous):

its not D.

hartnn (hartnn):

you need only 2 points, take any 2 of these 4 points and apply the formula

hartnn (hartnn):

The slope of the line through points (x1,y1) and (x2,y2) is given by : \(\huge m=\frac{y_1-y_2}{x_1-x_2}\) now,just put the values and find m.

hartnn (hartnn):

did u get slope as 1 ?

OpenStudy (anonymous):

I got -1, i probably did somthing wrong, i suckk at math

hartnn (hartnn):

lets take 2 points as 0,0 and 1,1 :)

hartnn (hartnn):

y1 - y2 = numerator = 1-0 = 1 x1-x2 = denominator = 1-0 = 1 so, m = 1/1 = 1 :)

OpenStudy (anonymous):

Really thank you soo much, your really helpful compared to other people. wow(:

hartnn (hartnn):

hehe, thanks for the compliment :) you are really beautiful compared to other people. wow(: and welcome ^_^

OpenStudy (anonymous):

omg. thank you.<3

hartnn (hartnn):

<3